Chapter 2 Applying Excel (B) Help Saved 1 Required information This Excel worksheet relates to the Dickson Company example that is summarized in Exhibit 2-5 10 points 2. Change the total fixed manufacturing overhead cost for the Milling Department in Data area back to $390,000, keeping all of the other data the same as in the original example. Consider a new job, Job 408, with the following characteristics: eBook Print A B C References 1 Chapter 2: Applying Excel 3 Cost summary for Job 408 Department 4 Milling Assembly 5 4 Machine-hours 90 6 Direct labor-hours 20 7 Direct materials cost 640 $ 230 8 $ Direct labor cost 320 162
EXHIBIT 2-5 Dickson Company: An Example Using Multiple Predetermined Overhead Rates Step 1: Calculate the estimated total manufacturing overhead cost for each department. Miling Department Overhead Cost (Y: Assembly Department Overhead Cost (Y): =$390,000+$120 0 MHx 60,000 MHs) =$510,000 DLH x 80,000 DLHS) 5500,000 = $500.000 + $300,000 390,000 2.00 $800,000 Step 2: Calculate the predetermined overhead rate in each department. Milling Department Overhead Rate: Assembly Department Overhedad Rate: $510,000 = 60 000 machine-hours $800.0 =80000 dlrect labor-hours $10.00 per direct labor-hour = $8.50 per machine-hour Step 3: Calculate the amount overhead applied from both departments to Job 407. MiNing Department: Overhead Applied to Job 407 Assembly Department: Overhead Applied to Job 407 =$8.50 per MHx 90 MHs = $ 765 $10.00 per DLH x 20 DLHS = $200 Step 4: Calculate the total job cost for Job 407. Assembly $370 $280 $200 Milling $800 Total Direct materials $1,170 350 Direct labor 70 $765 Manufacturing overhead applied 965 $2,485 Total cost of Job 407 Step 5: Calculate the selling price for Job 407. Total cost of Job 407 $2,485.00 Markup ($2,485 x 75% ) 1,863.75 $4,348.75 Selling price of Job 407
In your worksheet, enter this new data in the cells for Job 407. 10 points What is the new selling price for Job 408? (Round your final answer to decimal places.) eBook Selling price for Job 408 Print References 3. Without changing the data for the job from requirement 2 above, what is the selling price for Job 408 if the total number of machine- hours in the Assembly Department increases from 3,000 machine-hours to 6,000 machine-hours? (Round your final answer to 2 decimal places.) Selling price for Job 408 4. Restore the total number of machine-hours in the Assembly Department to 3,000 machine-hours. And keep the job data the same as it was in Requirement 2. What is the selling price for Job 408 if the total number of direct labor-hours in the Assembly Department decreases from 80,000 direct labor-hours to 50,000 direct labor-hours? (Round your final answer to 2 decimal places.) Selling price for Job 408
Answer #1

Ans-2 Total manufacturing Overhead cost Milling Department Assembly Department -500000+(3.75x80.000 DLHS) = 500000+300000 = $Ans-3 Total manufacturing Overhead cost Milling Department Assembly Department =500000+(3.75X80.000 DLHS) = 500000+300000 = $Ans-4 Total manufacturing Overhead cost Milling Department Assembly Department =500000+(3.75X50000 DLHs) 500000+187500 =$ 687